OpenStudy (anonymous):

In triangle ABC, what is the measure of c if a = 12, b = 25, and C = 20°? 14.32 22.07 23.74 205.18

Nnesha (nnesha):

to find c side \[\huge\rm c^2 = a^2 +b^2 -2ab \times \cos(c)\] replace by their values and then solve !

OpenStudy (anonymous):

12^2+25^2-2(12)(25)xcos?

Nnesha (nnesha):

cos C = ??

OpenStudy (anonymous):

20 de

Nnesha (nnesha):

yes right so replace C by 20 de

OpenStudy (anonymous):

Sides: a = 12   b = 25   c = 14.324 (so the first answer is correct) cos c ---> may be cos gamma http://www.triangle-calculator.com/?what=&q=a+%3D+12%2C+b+%3D+25%2C+and+C+%3D+20%C2%B0&submit=Solve

OpenStudy (anonymous):

I did get 14.324

OpenStudy (anonymous):

though

Nnesha (nnesha):

are you sure ?

OpenStudy (anonymous):

yea

Nnesha (nnesha):

i got different answer

OpenStudy (anonymous):

I hope im right

OpenStudy (anonymous):

what did you get ?

Nnesha (nnesha):

yes that's not right try again

OpenStudy (anonymous):

careful about degree / radian stuff!

OpenStudy (anonymous):

got 205

Nnesha (nnesha):

^ :-)

Nnesha (nnesha):

:-) = that's right